OpenStudy (anonymous):

27. What are the excluded values of the function? 7/2x + 48 (1 point) A.0.04 B.24 C.–24 D.–0.04

OpenStudy (anonymous):



OpenStudy (kropot72):

What would happen if the value of the denominator became zero?

OpenStudy (kropot72):

\[7/ 0=\infty\]

OpenStudy (anonymous):

i came up with D?

OpenStudy (kropot72):

Therefore the excluded value is found by putting the denominator equal to zero and solving as follows: 2x + 48 = 0 2x = -48 x = -24 the correct choice is c